OpenStudy (anonymous):

find y-int of f(x)=-5(0.7)

OpenStudy (anonymous):

whr r u from

OpenStudy (anonymous):

why does that matter

OpenStudy (anonymous):

yessirr

OpenStudy (anonymous):

gujurat

OpenStudy (anonymous):

ohh im from mumbai

OpenStudy (anonymous):

cool

OpenStudy (anonymous):

wat do u do

OpenStudy (anonymous):

school

OpenStudy (anonymous):

i think u have to multipy 7 and 5

OpenStudy (anonymous):

thanks ill try that!

OpenStudy (anonymous):

your missing a variable

OpenStudy (anonymous):

i think u have to find thex mean idk

OpenStudy (anonymous):

yeah the question is improper

OpenStudy (anonymous):

shizz i am -5(0.7)^ x

OpenStudy (anonymous):

or u have to divide no multiply idk

OpenStudy (anonymous):

f(x) = -5(0.7)^ x all the answer choices are coordinates

OpenStudy (anonymous):

is the question \[f(x) = -5(0.7)^{x}\]?

OpenStudy (anonymous):

yess

OpenStudy (anonymous):

okay to find y intercept, equate f(x) to 0. so 0 = -5(0.7)^x or, 0.7^x =0 this means that x is undefined or negative infinity, depending on how you look at it.

OpenStudy (anonymous):

sorry positive infinity

OpenStudy (anonymous):

thank you so much that makes sense

OpenStudy (anonymous):

im soo confuse

OpenStudy (anonymous):

you are welcome

OpenStudy (anonymous):

its an exponential decay.. alice smith. it wont cross the x-axis.

OpenStudy (anonymous):

the graph meets the x axis at positive infinity